Background pertinent to this problem is available in Interactive LearningWare 18.3. A uniform electric field exists everywhere in the x, y plane. This electric field has a magnitude of 3500 N/C and is directed in the positive x direction. A point charge -5.9 × 10-9 C is placed at the origin. Find the magnitude of the net electric field at (a) x = -0.19 m, (b) x = +0.19 m, and (c) y = +0.19 m.
Answer #1

A)

Electric field, E = 3500 + (9 x 10^9 x 5.9 x 10^-9)/0.19^2

E = 4970.91 N/C

B)

E = 3500 - (9 x 10^9 x 5.9 x 10^-9)/0.19^2

E = 2029.09 N/C

C)

E = 3500i + [(9 x 10^9 x 5.9 x 10^-9)/0.19^2] j

E = sqrt(3500^2 + 1479.91^2) = 3796.52 N/C

Comment in case any doubt please rate my answer....